# Branching Fixed Effects: A Proposal for Communicating Uncertainty

### v1

@@ Abstract
Economists often rely on estimates of linear fixed effects models developed by other teams of researchers. Assessing the uncertainty in these estimates can be challenging. I propose a form of sample splitting for network data that breaks two-way fixed effects estimates into statistically independent branches, each of which provides an unbiased estimate of the parameters of interest. These branches facilitate uncertainty quantification, moment estimation, and shrinkage. Algorithms are developed for efficiently extracting branches from large datasets. I illustrate these techniques using a benchmark dataset from Veneto, Italy that has been widely used to study firm wage effects.
@@ 1 Introduction
This paper proposes breaking two-way fixed effects estimates into statistically independent branches as a means of transparently communicating uncertainty. Each branch corresponds to a distinct subsample of the microdata within which unbiased estimates of all target parameters can be constructed. Building on a graph-theoretic interpretation of two-way fixed effects models developed in [kline2024firm], the branches are constructed from edge-disjoint spanning trees of the mobility network—a graph representing the evolution of group memberships in the data. Though I will focus on a setting where this network captures the structure of worker moves between firms, the spanning tree construction applies broadly to contexts involving two-dimensional heterogeneity, including student–teacher, patient–doctor, and judge–district pairings. Edges not included in any spanning tree are appended to the final branch, which ensures the full-sample fixed effects estimator decomposes linearly into branch-specific estimates.
Publicly releasing branches allows outside researchers to transparently assess uncertainty in published fixed effect estimates and in projections of those estimates onto observable covariates. Branches also simplify more advanced tasks such as estimating moments of the fixed effects and shrinkage that have proved challenging to extend to high-dimensional fixed effects settings with heteroscedasticity [kwon2023optimal, cheng2025optimal, robin2025]. I illustrate these ideas using as an example the estimation of firm fixed effects in a benchmark dataset from the Italian province of Veneto.
@@
The Veneto data provide a challenging test case for the branching approach, as many firms are connected by a single worker move. In this low connectivity environment, constructing even two branches requires limiting attention to 29% of the firms in the largest connected set. However, the estimates turn out to be remarkably insensitive to further pruning of the sample despite the fact that larger firms tend to be better connected. Among other interesting findings, the analysis reveals that the distribution of firm wage effects in these pruned samples are skew left and exhibit heavy tails. In other empirical literatures where the relevant graphs tend to exhibit higher edge-connectivity—e.g., empirical work on teacher value added or models of place effects—it may be feasible to extract dozens of branches without meaningfully narrowing the scope of investigation.
@@ 2 The blessings of branches
Since the branches are comprised of disjoint samples, the full-sample least squares estimator can be decomposed as the sum
@@
It is also useful to construct leave-out estimators $\left\{\hat{\psi}_{-b}\right\}_{b=1}^{M}$, where each $\hat{\psi}_{-b}$ gives the least squares estimator fit to the microdata in all branches but $b$. By assumption $\mathbb{E}\left[\hat{\psi}_{-b}\right]=\psi$. Therefore, a statistically independent estimator of $\phi_{b}$ is given by
$\hat{\phi}_{-b}=\boldsymbol{C}_{b}\hat{\psi}_{-b}.$
As detailed in Section 5, after the data have been partitioned into branches, computing each $\hat{\phi}_{b}$ and $\hat{\phi}_{-b}$ is no more difficult than computing $\hat{\psi}$.
Note that $\hat{\psi}_{b}$, $\hat{\phi}_{b}$, and $\hat{\phi}_{-b}$ are each $J\times 1$ vectors that can be stored as columns in a spreadsheet alongside the full-sample estimate $\hat{\psi}$. By the decomposition property in (1), releasing the full-sample estimate is redundant once columns corresponding to the $\left\{\hat{\phi}_{b}\right\}_{b=1}^{M}$ vectors have been provided.
Having outlined what branches are, I now preview how access to the branch-specific vectors $\left\{\hat{\psi}_{b},\hat{\phi}_{b},\hat{\phi}_{-b}\right\}_{b=1}^{M}$ can facilitate three common empirical tasks:
@@ 1. Quantifying uncertainty
$\mathbb{E}\left[\hat{\phi}_{b}\left(\hat{\phi}_{b}-\hat{\phi}_{-b}\right)^{\prime}\right]$ $=$ $\mathbb{E}\left[\left(\hat{\phi}_{b}-\hat{\phi}_{-b}\right)\hat{\phi}_{b}^{\prime}\right]$
$=$ $\mathbb{E}\left[\hat{\phi}_{b}\hat{\phi}_{b}^{\prime}\right]-\mathbb{E}\left[\hat{\phi}_{-b}\hat{\phi}_{b}^{\prime}\right]$
@@ 3. Shrinkage
Standard empirical bayes shrinkage arguments are predicated on the assumption that the distribution of noise is known ex-ante [waltersEB2024]. A recent paper by [ignatiadis2023empirical] proposes a best predictor based on independent and identically distributed replicates that adapts to the unknown noise distribution. I show that important insights from this paper carry over to the problem of predicting the elements of $\psi$ given independent branch estimates $\left\{\hat{\psi}_{b}\right\}_{b=1}^{M}$ that are not identically distributed. The proposed procedure consists of running a series of regressions of each entry of $\hat{\psi}_{b}$ against the values of that entry in the other branches. This process is repeated for each choice of $b$. Finally, the predicted values, which generally shrink the noisy branch-specific estimates, are averaged.
@@ 6 Building branches
In the prototypical worker–firm dataset, the edge connectivity is zero: there are usually multiple connected components of firms. Since the work of [abowd2002computing], the convention in the literature has been to focus on the largest connected component of the mobility graph (i.e., the largest $1$-ECC). Typically many of firms in the largest $1$-ECC are connected by only a single edge, in which case no more than one edge-disjoint spanning tree can be packed into the graph. The Nash-Williams–Tutte theorem suggests a reasonable way to find multiple disjoint spanning trees is to narrow the scope of investigation from the largest $1$-ECC to the largest $k$-ECC for $k>1$. Doing so requires first finding the largest $k$-ECC and then packing this subgraph.
@@ 6.2 Packing a $k$-ECC
${\max_{x}}$ $\sum_{T\in\mathcal{T}}x_{T}$
s.t. $\sum_{T\ni e}x_{T}\;\leq 1$ $\forall\,e\in E,$
$x_{T}\in\{0,1\}$ $\forall\,T\in\mathcal{T},$
@@ 7 Empirical application
I now turn to analyzing an extract of the benchmark Veneto dataset that was also studied in [kline2020leave] and [kline2024firm]. The extract is comprised of 1,859,459 person-year observations from the years 1999 and 2001. The largest connected component contains 73,933 firms and 747,205 workers, 197,572 of whom switch employers between the two years. The worker moves between these two years involve 150,417 ordered origin-destination pairs $\left|\mathcal{P}\right|$. Exactly 1,500 pairs of firms have flows in both directions. Hence, the mobility network formed by these pairs contains $\left|E\right|=$150,417-1,500=148,917 undirected edges.
@@
Comparing these findings to Table 1 reveals that randomly splitting workers into $M$ groups yields more firms with $M$ estimates than does pruning to the largest $k$-ECC with $M$ trees. For example, the largest $3$-ECC has 21,565 firms, while randomly assigning workers to one of two splits yields a median of 28,737 firms with two independent estimates. This discrepancy suggests that there exist subgraphs of the largest $2$-ECC that strictly nest the largest $3$-ECC yet contain two edge-disjoint spanning trees. How to systematically explore the space of such intermediate subgraphs is an interesting question. For example, one could potentially append pruned edges to the branches of the $3$-ECC, growing the size of the network in way that ensures each branch connects the same set of vertices. I leave such investigations to future research.
While random splitting delivers independent estimates for more firms, a major advantage of the pruning and packing strategy is that the target population is deterministic. Another advantage is that the full sample estimator decomposes linearly into $M$ branch-specific estimates, which facilitates uncertainty quantification. In contrast, the random splitting approach yields a full-sample estimator that includes a fixed effect for every firm in the union of the $M$ connected sets. While some of these parameters have $M$ independent estimates, others have may have only a single estimate, which hinders uncertainty quantification even when ignoring the randomness in the split itself.
@@ 7.2 Moment estimation
Table 5 converts the average moment estimates from Table 4 into scaled moments, which facilitates comparison to a Gaussian benchmark. The scaled higher moment estimates suggest that the size-weighted distribution of firm effects exhibits a left skew: more workers are employed at firms with very low than very high wages. This negative skew is less pronounced in the $6$-ECC, perhaps because firms in the $6$-ECC tend to be very large.
@@ 7.3 Shrinkage
where the second equality follows by independence of the branch measurement errors. Thus, a regression of the estimates in one branch on those from another yields a minimum mean squared error optimal predictor of the latent firm effect regardless of the noise distributions $\left\{F_{1},F_{2}\right\}$. Importantly, this result holds under arbitrary patterns of heteroscedasticity, both across and within branches.

### v2

@@ Abstract
Economists often rely on estimates of linear fixed effects models produced by other teams of researchers. Assessing the uncertainty in these estimates can be challenging. I propose a form of sample splitting for networks that partitions the data into statistically independent branches, each of which can be used to compute an unbiased estimate of the parameters of interest in two-way fixed effects models. These branches facilitate uncertainty quantification, moment estimation, and shrinkage. Drawing on results from the graph theory literature on tree packing, I develop algorithms to efficiently extract branches from large networks. I illustrate these techniques using a benchmark dataset from Veneto, Italy that has been widely used to study firm wage effects.
@@ 1 Introduction
This paper proposes breaking two-way fixed effects estimates into statistically independent branches as a means of transparently communicating uncertainty. Each branch corresponds to a distinct subsample of the microdata within which unbiased estimates of all target parameters can be constructed. Building on a graph-theoretic interpretation of two-way fixed effects models developed in [kline2024firm], the branches are constructed from edge-disjoint spanning trees of the mobility network—a graph representing the evolution of group memberships in the data. Though I will focus on a setting where this network captures the structure of worker moves between firms, the spanning tree construction applies broadly to contexts involving two-dimensional heterogeneity, including student–teacher, patient–doctor, and judge–district pairings. Edges not included in any spanning tree are appended to the last tree’s branch, which ensures the full-sample fixed effects estimator decomposes linearly into branch-specific estimates.
Publicly releasing branches allows outside researchers to transparently assess uncertainty in published fixed effect estimates and in projections of those estimates onto observable covariates. Branches also simplify more advanced tasks such as estimating moments of the fixed effects and shrinkage that have proved challenging to extend to high-dimensional fixed effects settings with heteroscedasticity [kwon2023optimal, cheng2025optimal, robin2025]. I illustrate these ideas using as an example the estimation of firm wage fixed effects in a benchmark dataset from the Italian province of Veneto.
@@
The Veneto data provide a challenging test case for the branching approach, as many firms are connected by a single worker move. In this low connectivity environment, constructing even two branches requires limiting attention to 29% of the firms in the largest connected set. However, the estimates turn out to be remarkably insensitive to further pruning of the sample despite the fact that larger firms tend to be better connected. Among other interesting findings, the analysis reveals that the firm wage fixed effects in these pruned samples are left-skewed and heavy-tailed. In other empirical literatures where the relevant graphs tend to exhibit higher edge-connectivity—e.g., empirical work on teacher value added or models of place effects—it may be feasible to extract dozens of branches without meaningfully narrowing the scope of investigation.
@@ 2 The blessings of branches
Since the branches correspond to disjoint subsamples, the full-sample least squares estimator can be decomposed as the sum
@@
It is also useful to construct leave-out estimators $\left\{\hat{\psi}_{-b}\right\}_{b=1}^{M}$, where each $\hat{\psi}_{-b}$ gives the least squares estimator fit to the microdata in all branches but $b$. By assumption $\mathbb{E}\left[\hat{\psi}_{-b}\right]=\psi$. Therefore, a statistically independent estimator of $\phi_{b}$ is given by $\hat{\phi}_{-b}=\boldsymbol{C}_{b}\hat{\psi}_{-b}.$ As detailed in Section 5, after the data have been partitioned into branches, computing each $\hat{\phi}_{b}$ and $\hat{\phi}_{-b}$ is no more difficult than computing $\hat{\psi}$.
My proposal is for researchers to share the branch-level vectors $\left\{\hat{\psi}_{b},\hat{\phi}_{b},\hat{\phi}_{-b}\right\}_{b=1}^{M}$ with the public. Note that $\hat{\psi}_{b}$, $\hat{\phi}_{b}$, and $\hat{\phi}_{-b}$ are each $J\times 1$ vectors that can be stored as columns in a spreadsheet alongside the full-sample estimate $\hat{\psi}$. By the decomposition property in (1), releasing the full-sample estimate is redundant once columns corresponding to the $\left\{\hat{\phi}_{b}\right\}_{b=1}^{M}$ vectors have been provided.
Having outlined what branches are, I now preview how access to $\left\{\hat{\psi}_{b},\hat{\phi}_{b},\hat{\phi}_{-b}\right\}_{b=1}^{M}$ can facilitate three common empirical tasks:
@@ 1. Quantifying uncertainty
$\mathbb{E}\left[\hat{\phi}_{b}\left(\hat{\phi}_{b}-\hat{\phi}_{-b}\right)^{\prime}\right]$ $=$ $\mathbb{E}\left[\hat{\phi}_{b}\hat{\phi}_{b}^{\prime}\right]-\mathbb{E}\left[\hat{\phi}_{b}\hat{\phi}_{-b}^{\prime}\right]$
@@ 3. Shrinkage
Standard empirical Bayes shrinkage arguments are predicated on the assumption that the distribution of noise is known ex-ante [waltersEB2024]. A recent paper by [ignatiadis2023empirical] proposes a best predictor based on independent and identically distributed replicates that adapts to the unknown noise distribution. Insights from this paper carry over to the problem of predicting the elements of $\psi$ given independent branch estimates $\left\{\hat{\psi}_{b}\right\}_{b=1}^{M}$ that are not identically distributed. The proposed procedure, which builds on a suggestion by [krutchkoff1967supplementary], consists of running a series of regressions of each entry of $\hat{\psi}_{b}$ against the values of that entry in the other branches. This process is repeated for each choice of $b$. Finally, the predicted values, which generally shrink the noisy branch-specific estimates, are averaged.
@@ 6 Building branches
In the prototypical worker–firm dataset, the edge connectivity is zero: there are usually multiple connected components of firms. Since the work of [abowd2002computing], the convention in the literature has been to focus on the largest connected component of the mobility graph (i.e., the largest $1$-ECC). Typically many firms in the largest $1$-ECC are connected by only a single edge, in which case no more than one edge-disjoint spanning tree can be packed into the graph. The Nash-Williams–Tutte theorem suggests a reasonable way to find multiple disjoint spanning trees is to narrow the scope of investigation from the largest $1$-ECC to the largest $k$-ECC for $k>1$. Doing so requires first finding the largest $k$-ECC and then packing this subgraph.
@@ 6.2 Packing a $k$-ECC
$\max_{x}\ \sum_{T\in\mathcal{T}}x_{T}\quad\text{s.t.}\quad\sum_{T\ni e}x_{T}\leq 1,\ \forall e\in E,\qquad x_{T}\in\{0,1\},\ \forall T\in\mathcal{T},$
@@ 7 Empirical application
I now turn to analyzing an extract of the benchmark Veneto dataset that was also studied in [kline2020leave] and [kline2024firm]. The extract consists of 1,859,459 person-year observations from the years 1999 and 2001. The largest connected component contains 73,933 firms and 747,205 workers, 197,572 of whom switch employers between the two years. The worker moves between these two years involve 150,417 ordered origin-destination pairs $\left|\mathcal{P}\right|$. Exactly 1,500 pairs of firms have flows in both directions. Hence, the mobility network formed by these pairs contains $\left|E\right|=$150,417-1,500=148,917 undirected edges.
@@
Comparing these findings to Table 1 reveals that randomly splitting workers into $M$ groups yields more firms with $M$ estimates than does pruning to the largest $k$-ECC with $M$ trees. For example, the largest $3$-ECC has 21,565 firms, while randomly assigning workers to one of two splits yields a median of 28,737 firms with two independent estimates. This discrepancy suggests that there exist subgraphs of the largest $2$-ECC that strictly nest the largest $3$-ECC yet contain two edge-disjoint spanning trees. How to systematically explore the space of such intermediate subgraphs is an interesting question. For example, one could potentially append pruned edges to the branches of the $3$-ECC, growing the size of the network in a way that ensures each branch connects the same set of vertices. I leave such investigations to future research.
While random splitting delivers independent estimates for more firms, a major advantage of the pruning and packing strategy is that the target population is deterministic. Another advantage is that the full sample estimator decomposes linearly into $M$ branch-specific estimates, which facilitates uncertainty quantification. In contrast, the random splitting approach yields a full-sample estimator that includes a fixed effect for every firm in the union of the $M$ connected sets. While some of these parameters have $M$ independent estimates, others may have only a single estimate, which hinders uncertainty quantification even when ignoring the randomness in the split itself.
@@ 7.2 Moment estimation
Table 5 converts the average moment estimates from Table 4 into scaled moments, which facilitates comparison to a Gaussian benchmark. The scaled higher moment estimates suggest that the size-weighted distribution of firm effects exhibits a left skew: more workers are employed at firms with very low wages than at firms with very high wages. This negative skew is less pronounced in the $6$-ECC, perhaps because firms in the $6$-ECC tend to be very large.
@@ 7.3 Shrinkage
where the second equality follows by independence of the branch measurement errors. Thus, a regression of the estimates in one branch on those from another yields a minimum mean squared error optimal predictor of the latent firm effect regardless of the noise distributions $\left\{F_{1},F_{2}\right\}$, an insight that can be traced back at least to [krutchkoff1967supplementary]. Importantly, this result holds under arbitrary patterns of heteroscedasticity, both across and within branches.